OpenStudy (anonymous):

Which of the following shows the square completed in the this equation? x2 + 3x - 5 = 0 A. (x + 9)2 = 14 B. (x - 3)2 = 8 C. (x + 1.5)2 = 7.25

OpenStudy (vishal_kothari):

x^2 + 3x - 5 = 0 x^2 + 3x + (3/2)^2 - 5 = (3/2)^2 (x + 3/2)^2 = 9/4 + 5 (x + 1.5^)2 = 7.25 So the answer is C..

OpenStudy (anonymous):

thnk u vishal

OpenStudy (vishal_kothari):

welcome...

OpenStudy (anonymous):